OpenStudy (anonymous):

sin2 + (.6)^2 = 1 sin2 = .64 sin = .8 how is the 6^2 changed to .64

OpenStudy (jhannybean):

\[\large \sin^2(x) +(0.6)^2 = 1\]\[\large \sin^2(x) = 1-(0.6)^2\]\[\large \sin(x) = \sqrt{1-(0.6)^2}\]\[\large \sin(x) = 0.8\] That's how it changed.